OpenStudy (wesdg1978):

Given that f(3)=-1 and f'(3)=5, find an equation for the tangent line to the graph of y=f(x) at x=3 How do I do this?

OpenStudy (anonymous):

use the point slope formula \[y-y_1=m(x-x_1)\] with \[x_1=3, y_1=-1, m=5\]

OpenStudy (wesdg1978):

thanks